OpenStudy (anonymous):

Solve (cubed root of x - 3) – 5 = 16 Answer x = 5 x = 19 x = 52 x = 67

OpenStudy (anonymous):

@legoauto220

OpenStudy (maheshmeghwal9):

\[(\sqrt[3]{x}-3)-5=16.\]Is this ur question?

OpenStudy (anonymous):

x-3 is under the cubed root

OpenStudy (anonymous):

\[\sqrt[3]{x-3}-5=16\]

OpenStudy (maheshmeghwal9):

k:) \[\sqrt[3]{x-3}=16+5=21.\]\[\implies (x-3)=21^3.\]\[\implies x=21^3+3.\]

OpenStudy (maheshmeghwal9):

just calculate now:)

OpenStudy (anonymous):

i know i did that but thats not one of the options above

OpenStudy (anonymous):

\[\huge \sqrt[3]{x-3}-5=16\] \[\huge \sqrt[3]{x-3}=21\] \[\huge x-3=9261\] \[\huge x=9264\] Can't be what you meant...

OpenStudy (maheshmeghwal9):

I did the same options are wrong:)

OpenStudy (anonymous):

i guess its just 3 times the square root i got the answer now 52

OpenStudy (anonymous):

x^3 = x cubed x\(^{1/3}\) = \(\sqrt[3]{x}\) = cube root of x

OpenStudy (anonymous):

The are NOT the same thing, at all :D

OpenStudy (anonymous):

i know i wish i could show you how its displayed because it shows the cubed root

OpenStudy (maheshmeghwal9):

ya @cplatt688 that must be if u want answer:) but u gt 2 question's solution one with cuberoot & one with 3 x......

OpenStudy (anonymous):

3 sqrt (x) = 3x^(1/2)

OpenStudy (anonymous):

Or in clearer writing: \[3\sqrt{x} = 3x^{1/2}\]

OpenStudy (anonymous):

Got to always remember to be extra, extra careful with exponents (or fractional exponents, which are better known as "roots")